If you throw a snowball straight upwards with a velocity of 19 m/s, what will the magnitude of its velocity be after 2.9 seconds?
1
Expert's answer
2018-02-18T09:10:07-0500
v = v0 - gt = 19 - 9.81 x 2.9 = -9.45 m/s
The magnitude will be 9.45 m/s.